Using the Put-Call Parity relationship, find the value of a put option (same strike, expiration as the call option) using the following information:
Current stock price $33 risk-free rate 6%
Call option strike price $32 Call option current value $6.56
Option time to maturity 1 years
Answer #1

According to call put parity theorem ,

Call price + [Strike price *e^-rt] = Put price +current price

6.56 + [32*e^-.06*1] = P + 33

6.56 + [32 * e^-.06] = P +33

6.56 + [32*.94176] = P +33

6.56 + 30.13632 = P +33

P = 36.69632 - 33

= 3.69632 (rounded to 3.70 )

**Find value of e ^-.06 using financial calculator